OpenStudy (anonymous):

simplify radical 18- radical 2

OpenStudy (anonymous):

\[\sqrt{18}-\sqrt{2}=\sqrt{3*3*2}-\sqrt{2}\]

OpenStudy (anonymous):

rad2(3-1)

OpenStudy (anonymous):

2rad2

OpenStudy (anonymous):

\[3\sqrt{2}-\sqrt{2}=\sqrt{2}\]

OpenStudy (anonymous):

Ummm...@nancy? 3rad(2) - rad(2) = rad(2)? I think you mean\[3\sqrt{2}-\sqrt{2}=2\sqrt{2}\]